If A1​,…,An are Bernoulli random variables with parameter 0.4 and E[A1+...+An]=8.4, then what is n?
Answer #1

E[4] + A2 + ... + An] =8.4 =E(4) +E(A2)+...+E(An) = 8.4 0.4+0.4+... (n times) = 8.4 0.4n = 8.4 =n=84 = 21 0.4